# Rad4XCNN: a new agnostic method for post-hoc global explanation of CNN-derived features by means of radiomics

### v1

@@ 4.2 Traditional CNN Explainability
As the most accurate model in the external test, ResNet was selected to analyze the traditional explainability methods based on saliency maps. By comparing the saliency maps obtained with GradCAM, EigenCAM, and ScoreCAM methods, it is possible to observe conflicting results among these interpretability techniques. Figure 2 shows some examples. Despite ResNet’s robust performance, the interpretations provided by these methods often diverged, presenting a challenge in understanding the model decision. This discrepancy raises questions about the reliability and consistency of these interpretability methods, urging a closer examination of their underlying mechanisms. Similar inconsistencies are recently achieved in Cerekci et al. (2024); Zhang et al. (2021). In particular, rows 1-2 show two examples of correctly classified lesions, while Rows 3-4 show two examples of wrongly classified lesions. Some explanations predominantly focused on background regions, even when the model’s prediction was correct. Other concerns arise in the case of wrong prediction. Considering Grad-CAM, this effect is due to the dependence of the gradient calculation of $\frac{\partial y^{c}}{\partial A^{k}_{ij}}$ (see Equation 2) that is a function of the prediction $y^{c}$, assumed correctly predicted the class $c$. For Score-CAM this problem is mitigated by CIC which weights the importance of each activation map. For Eigen-CAM decomposition does not depend on the predicted class and gradient computation. This phenomenon underlines the complexity of interpreting neural network decisions and highlights the need for newer interpretability techniques capable of overcoming these problems and providing clearer and more actionable insights.
@@
benign lesion
(correctly classified)
@@
malignant lesion
(correctly classified)
@@
benign lesion
(wrongly classified)
@@
malignant lesion
(wrongly classified)
@@
4.3 Explaining Deep Features by means of Radiomics in US Breast Dataset
Figure 3 depicts the Rad4XCNN results, utilizing four distinct threshold values $\rho=[0.30,0.35,0.40,0.45]$. Notably, the graph exclusively illustrates correlations with deep features derived from convolutional networks, as no correlations were identified with features extracted from VITs. Regarding the CNN-derived features, our analysis reveals that Energy, and TotalEnergy emerge as the most prominently correlated features, with 15 correlations observed at a threshold of 0.3, 9 at 0.35, 4 at 0.4, and 2 at 0.45. Additionally, a high correlation (with $\rho=0.45$) was established with SizeZoneNonUniformity, DependenceNonUniformity, and RunLengthNonUniformity. These 5 radiomic features (i.e. Energy, TotalEnergy, SizeZoneNonUniformity, DependenceNonUniformity, and RunLengthNonUniformity) collect all the correlations, both those with the original (unfiltered) version of the feature and those derived after filtering (considering the wavelets, with the 4 decompositions).
@@ 5.1 Comparing ViT-derived, CNN-derived and Radiomic Features
The ability of deep architectures to extract high and low level features makes deep features much more informative than radiomic ones. As a consequence, many works have shown that approaches exploiting deep features achieve higher performance than radiomics Lisson et al. (2022); Sun et al. (2020); Truhn et al. (2019); Wei (2021). Hence, considering the evaluation on the external test, CNN-derived features enabled higher accuracy compared with ViT-derived. Regardless of datasets, ResNet maintains high accuracy values.
In terms of explainability
@@
Saliency maps generated through methods like GradCAM, EigenCAM, and ScoreCAM can effectively emphasize meaningful areas for prediction. While they serve as valuable visual and qualitative aids for prediction analysis, they present two primary concerns: i) each method yields distinct explanations, indicating a need for cautious interpretation; ii) these methods can lead to inter-operator variability evaluation, as individual physicians may interpret the explanations differently; iii) these explanations are inherently local, posing challenges for clinical validation of the models. In contrast, radiomic features offer clearer and more comparable explanations. Since it is possible to associate each radiomic feature with meaning/behavior related to its analytical definition, they enable the comparison of features identified in other studies with those observed in clinical practice, thereby enhancing their utility and reducing ambiguity.
Comparing CNN-derived and ViT-derived features
@@
As highlighted in Section 4.3, a notable alignment was observed between features derived from convolutional networks, whereas none was found with those extracted from ViT. This outcome aligns intuitively with an examination of the internal structures of the two architectures. CNN-based approaches typically rely on hierarchical feature extraction through a series of convolutional and pooling layers, capturing local patterns and spatial dependencies within the image. In contrast, ViT architectures adopt a self-attention mechanism, enabling them to capture global context by considering relationships between all image patches simultaneously. This leads CNNs to excel at capturing fine-grained details and local structures while ViTs in modeling long-range dependencies and contextual understanding. The authors themselves acknowledge that Transformers lack certain inherent biases found in convolutional neural networks, such as translation equivariance and locality Dosovitskiy et al. (2020). This similarity between radiomic features and CNN-derived features arises from their shared principle of extracting local characteristics. Convolutional filters in CNNs capture local information through the Receptive Field kernels and a hierarchical structure of layers. Likewise, radiomic features are predominantly derived from matrices that analyze the relationships among gray levels of adjacent pixels. This common approach aligns CNN-derived features closely with radiomic features, setting them apart from ViT-derived features. For these reasons, Rad4XCNN is a global explanation method for CNN-derived features.
@@ 5.2 Mitigating XAI Concerns Through Rad4XCNN
In our research, we uncovered a significant inconsistency in methods relying on saliency maps, a finding corroborated by numerous other studies Cerekci et al. (2024); Prinzi et al. (2024a); Zhang et al. (2021). Consequently, even though these methods can identify where a model focuses for prediction, their reliability remains doubtful Cerekci et al. (2024); Oh et al. (2020); Prinzi et al. (2023a, 2024a); Signoroni et al. (2021); Zhang et al. (2021). As a result, it is impossible to state with certainty the reliability of the trained model and to extract certain clinical insights.
@@
In contrast, Rad4XCNN presents a novel approach distinct from visualization map-based methods, which is a new challenge in the realm of explainable methods in medical image analysis Papanastasiou et al. (2023). This approach allows medical professionals to analyze whether radiomic features, whose significance is well defined, are in line with the clinical literature. In addition, it is possible to extract new clinical findings, such as those discussed in Section 5.3. This aspect overcomes the human tendency towards a positive interpretation of results by choosing the solution they want to be the correct one Bornstein (2016), providing an objective explanation.
Accuracy and Explainability trade-off Dilemma
@@
Achieving the trade-off between explainability and accuracy is still an open dilemma Bell et al. (2022). Although this trade-off has been widely discussed, there are some experiments where it is highlighted that in the clinical context, a patient would prefer an accurate model rather than an explainable one van der Veer et al. (2021). This would suggest that one should never sacrifice accuracy for the sake of explainability. Considering the absence of a definitive solution to this question, Rad4XCNN refrains from imposing constraints on model training. Instead, it provides an approach where radiomic features, although less predictive than deep ones, are exclusively utilized for explanatory purposes. It remains a task of the deep model to perform prediction accurately.
Explaining apparently incomprehensible patterns
@@
The abstraction mechanism within convolutional networks implies that the extracted features have a significantly higher level of abstraction than radiomic features. In this perspective, XAI methodologies must recognize that although it is possible to explain model decisions, understanding the relationships identified by deep architectures may be inherently incomprehensible. However, Rad4XCNN elucidates that a correlation exists between CNN-derived and radiomic features, demonstrating a convergence between physician insights and AI-based interpretations, as highlighted in the following section.
@@ 5.3 Radiomics-based explanation and its clinical validation
Rad4XCNN identified the radiomic features that exhibit the highest correlation with CNN-derived features, enabling us to draw significant clinical conclusions. Many of the radiomic features we discuss have already been used in previous studies, both diagnostic and prognostic, on the US radiomic analysis of focal breast lesions. For example, in the study by Cui et al. Cui et al. (2023) on the link between US radiomic features and biological functions in the prediction of HER2 status in breast cancer, Size Zone Non-Uniformity (SZNU) feature was associated with 1,871 genes in 75 perturbations and with carbon metabolism in cancer regarding biological functions; another radiomic feature analyzed in the same study Cui et al. (2023) was Run Length Non-Uniformity (RLNU), associated with the cell cycle and intercellular communication. In the study by Youk et al. on the analysis of the US radiomic features in benign and malignant breast masses Youk et al. (2020), 22 grayscale radiomic features were selected, some of which are also used in our study, such as Kurtosis, Energy, Run Length non-uniformity. Other radiomic features, although not yet applied to US image analysis, have already been analyzed in Prinzi et al. on breast cancer classification in DCE-MRI sequences Prinzi et al. (2024c). The main results concerned features closely related to the intensity of grey levels: Energy and Total Energy. These radiomic features describe the change in grey intensity after administration of contrast medium, which is more rapid in malignant lesions, so these latter shows a more rapid increase in energy, viceversa for benign lesions. In breast US evaluation, excluding completely anechoic lesions, grey levels related to the echogenicity of a focal lesion may vary, but most malignant lesions are hypoechoic Rahbar et al. (1999). Although the correlation between grey level intensity and the malignant lesion is different in DCE-MRI than in ultrasonographic features, even in ultrasonographic radiomic analysis, we can assert that both Energy and Total Energy emerge as characterizing features in ultrasonography, thereby laying a solid foundation for subsequent clinical investigations. The significance of signal intensity in US has been demonstrated in various anatomical districts Papini et al. (2002); Tessler et al. (2017), indicating its potential importance for breast imaging as well.

### v2

@@ 4.2 Traditional CNN Explainability
As the most accurate model in the external validation, ResNet was selected to analyze the traditional explainability methods based on saliency maps. By comparing the saliency maps obtained with GradCAM, EigenCAM, and ScoreCAM methods, it is possible to observe conflicting results among these interpretability techniques. Figure 3 shows some examples. Despite ResNet’s robust performance, the interpretations provided by these methods often diverged, presenting a challenge in understanding the model decision. This discrepancy raises questions about the reliability and consistency of these interpretability methods, urging a closer examination of their underlying mechanisms. Similar inconsistencies are recently achieved in Cerekci et al. (2024); Zhang et al. (2022). In particular, Rows 1-2 show two examples of correctly classified lesions, while Rows 3-4 show two examples of wrongly classified lesions. Some explanations predominantly focused on background regions, even when the model’s prediction was correct. Other concerns arise in the case of wrong prediction. Considering Grad-CAM, this effect is due to the dependence of the gradient calculation of $\frac{\partial y^{c}}{\partial A^{k}_{ij}}$ (see Equation 2) that is a function of the prediction $y^{c}$, assumed correctly predicted the class $c$. For Score-CAM this problem is mitigated by CIC which weights the importance of each activation map. For Eigen-CAM decomposition does not depend on the predicted class and gradient computation. This phenomenon underlines the complexity of interpreting neural network decisions and highlights the need for newer interpretability techniques capable of overcoming these problems and providing clearer and more actionable insights.
@@
Row 1:
benign lesion (correctly classified)
@@
Row 2:
malignant lesion (correctly classified)
@@
Row 3:
benign lesion (wrongly classified)
@@
Row 4:
malignant lesion (wrongly classified)
@@
4.3 Explaining Deep Features by means of Radiomics in Ultrasound Breast Dataset
Figure 4 depicts the Rad4XCNN results, utilizing four distinct threshold values $\rho=[0.30,0.35,0.40,0.45]$. The graph exclusively illustrates correlations with deep features derived from convolutional networks, as no correlations were identified with features extracted from VITs. Regarding the CNN-derived features (extracted from ResNet), our analysis reveals that Energy, and TotalEnergy emerge as the most prominently correlated features, with 15 correlations observed at a threshold of 0.3, 9 at 0.35, 4 at 0.4, and 2 at 0.45. Additionally, a high correlation (with $\rho=0.45$) was established with SizeZoneNonUniformity, DependenceNonUniformity, and RunLengthNonUniformity. These 5 radiomic features (i.e. Energy, TotalEnergy, SizeZoneNonUniformity, DependenceNonUniformity, and RunLengthNonUniformity) collect all the correlations, both those with the original (unfiltered) version of the feature and those derived after filtering (considering the wavelets, with the 4 decompositions).
@@ 5.1 Comparing ViT-derived, CNN-derived and Radiomic Features
The ability of deep architectures to extract high and low-level features makes deep features much more informative than radiomic ones. As a consequence, many works have shown that approaches exploiting deep features achieve higher performance than radiomics Lisson et al. (2022); Sun et al. (2020); Truhn et al. (2019); Wei (2021). In our case, considering the evaluation on the external validation, CNN-derived features enabled higher accuracy compared with ViT-derived (5). Regardless of datasets, ResNet maintains high accuracy values.
In terms of explainability -
@@
Saliency maps generated through GradCAM, EigenCAM, and ScoreCAM can effectively emphasize meaningful areas for prediction. While they serve as valuable visual and qualitative aids for prediction analysis, two primary concerns are presented: i) each method yields distinct explanations, indicating a need for cautious interpretation; ii) these methods can lead to inter-operator variability evaluation, as individual physicians may interpret the explanations differently; iii) these explanations are inherently local, posing challenges to extract global insights. In contrast, radiomic features offer clearer and more comparable explanations. Since it is possible to associate each radiomic feature with meaning/behavior related to its analytical definition, they enable the comparison of features identified in other studies with those observed in clinical practice.
Comparing CNN-derived and ViT-derived features -
@@
As highlighted in Section 4.3, a notable alignment was observed between CNN-derived and radiomic features, whereas none was found with those extracted from ViT. This outcome aligns intuitively with an examination of the internal structures of the two architectures. CNN-based approaches typically rely on hierarchical feature extraction through a series of convolutional and pooling layers, capturing local patterns and spatial dependencies within the image. In contrast, ViT architectures adopt a self-attention mechanism, enabling them to capture global context by considering relationships between all image patches simultaneously. This leads CNNs to excel at capturing fine-grained details and local structures while ViTs in modeling long-range dependencies and contextual understanding. The authors explicitly admit that some intrinsic biases present in convolutional neural networks—such as translation equivariance and locality—are absent from transformers Dosovitskiy et al. (2020). This similarity between radiomic features and CNN-derived features arises from their shared principle of extracting local characteristics. Convolutional filters in CNNs capture local information through the Receptive Field kernels and a hierarchical structure of layers. Likewise, radiomic features are predominantly derived from matrices that analyze the relationships among gray levels of adjacent pixels. This common approach aligns CNN-derived features closely with radiomic features, setting them apart from ViT-derived features. For these reasons, we consider Rad4XCNN as a global explanation method to explain CNN-derived features.
5.2 Explainability Properties of Rad4XCNN
Rad4XCNN offers a distinct perspective by explaining deep features through associations with the intrinsic interpretable radiomic features. This approach enables a richer clinically relevant understanding that surpasses the limitations of activation map-based methods. The advantages of Rad4XCNN can be summarized as follows:
  • provides global explanations: it enables global insights, facilitating the extraction of clinically meaningful patterns across the entire dataset. This global perspective is valuable for clinical applications, where understanding overarching trends is as critical as individual predictions;
  • uses intelligible features: the interpretable nature of radiomic features allows radiologists to correlate them with established clinical and radiological knowledge. While activation maps highlight regions of interest and enable only a qualitative analysis, radiomic features offer quantitative insights, helping bridge model predictions with well-known clinical evidence;
  • maintains model accuracy without training constraints: approaches try to overcome the explainability-accuracy trade-off (e.g., Self-Explaining Neural Networks) often encounter a trade-off between robustness and accuracy due to imposed architectural constraints Alvarez Melis and Jaakkola (2018). In fact, a change in test accuracy was shown as robustness regularization increases Elbaghdadi et al. (2020). As a result, these model categories increase robustness and explainability at the cost of accuracy. Other methods require a specific structure for their use. For example, traditional activation-based techniques such as CAM Zhou et al. (2016b), necessitate specific architectural configurations, like a Global Average Pooling layer before dense layers, limiting architecture flexibility. In contrast, Rad4XCNN allows the optimization of any CNN model using a standard training process. Then, in a post-hoc manner, the correlation between deep and radiomic features is computed, thus preserving model performance and flexibility across tasks. For this reason, Rad4XCNN i) achieves transparency without compromising predictive accuracy, ii) mitigates some concerns related to the explainability-accuracy trade-off, and iii) overcomes the accuracy-explainability trade-off dilemma.
@@ 5.3 Mitigating XAI Concerns Through Rad4XCNN
In our research, the significant inconsistencies in methods relying on saliency maps were confirmed, a finding corroborated by numerous other studies Cerekci et al. (2024); Prinzi et al. (2024a); Zhang et al. (2022). Consequently, even though these methods can identify where a model focuses for prediction, their reliability remains doubtful Cerekci et al. (2024); Oh et al. (2020); Prinzi et al. (2023a, 2024a); Signoroni et al. (2021); Zhang et al. (2022). As a result, it is impossible to state with certainty the reliability of the trained model and to extract certain clinical insights. Furthermore, the need to propose new methods beyond the visualization map-based methods was highlighted. Papanastasiou et al. (2023).
In light of this, we conducted a qualitative comparison of Rad4XCNN with three widely used saliency map generation methods, to prove Rad4XCNN reliability (see Figure 3). The exploited techniques in the comparison are among the most representative and widespread in the literature for explaining deep architecture predictions: GradCAM, is representative of a gradient-based method, while ScoreCAM and EigenCAM avoid gradient computation to overcome some GradCAM shortcomings. In this perspective, our results highlight the unreliability of the saliency map methods and comply with previous papers Zhang et al. (2022); Prinzi et al. (2024a). In particular, regarding Figure 3 we can highlight two main concerns:
  • discrepancy: the three methods return different explanations. In particular, for the image in Row 1, all the explanations differ. For the other images (Rows 2, 3, and 4), GradCAM explanation is deeply different from ScoreCAM and EigenCAM. This can negatively affect physician interpretation;
  • unreliability: although the first two images (Rows 1 and 2) show correct predictions, explanations seem unreliable. In fact, saliency maps highlight the background as the most important part.
@@
Rad4XCNN allows medical professionals to analyze whether radiomic features, whose significance is well defined, are aligned with the clinical literature. In addition, it is possible to extract new clinical findings, such as those discussed in Section 5.4. This aspect overcomes the human tendency towards a positive interpretation of results by choosing the solution they want to be the correct one Bornstein (2016), providing an objective explanation.
Accuracy and Explainability trade-off Dilemma -
@@
Achieving the trade-off between explainability and accuracy is still an open dilemma Bell et al. (2022). Although this trade-off has been widely discussed, there are some experiments where it is highlighted that in the clinical context, a patient would prefer an accurate model rather than an explainable one van der Veer et al. (2021b). This would suggest that one should never sacrifice accuracy for the sake of explainability. Considering the absence of a definitive solution to this question, Rad4XCNN refrains from imposing constraints on model training. Instead, it provides an approach where radiomic features, although less predictive than deep ones, are exclusively utilized for explanatory purposes. It remains a task of the deep model to perform prediction accurately.
Explaining apparently incomprehensible patterns -
@@
The abstraction mechanism within convolutional networks implies that the extracted features have a significantly higher level of abstraction than radiomic features. In this perspective, XAI methodologies must recognize that, although it is possible to explain model decisions, understanding the relationships identified by deep architectures may be inherently incomprehensible. However, Rad4XCNN elucidates that a correlation exists between CNN-derived and radiomic features, demonstrating a convergence between physician insights and AI-based interpretations, as highlighted in the following section.
@@ 5.4 Radiomics-based explanation and its clinical validation
Figure 5: Correlation trend between radiomic and deep features.
@@
The radiomic features, found to be correlated with the CNN-derived features (Figure 4), have already been used in literature breast cancer studies, both diagnostic and prognostic. It is worth noting that if a study identifies a specific feature as important, it allows for direct comparisons to analyze if the same feature holds significance in other studies. This is not feasible with only qualitative saliency maps. Radiomic features, representing and quantifying the manifestation of the underlying tumor phenotype, can be considered biomarkers to be integrated into clinical practice. As a result, they offer a quantitative basis for influencing clinical decision-making. For example, in the study by Cui et al. Cui et al. (2023) on the link between ultrasound radiomic features and biological functions in the prediction of HER2 status in breast cancer, Size Zone Non-Uniformity (SZNU) feature was associated with 1,871 genes in 75 perturbations and with carbon metabolism in cancer regarding biological functions; another radiomic feature analyzed in the same study Cui et al. (2023) was Run Length Non-Uniformity (RLNU), associated with the cell cycle and intercellular communication. In the study by Youk et al. on the analysis of the ultrasound radiomic features in benign and malignant breast masses Youk et al. (2020), 22 grayscale radiomic features were selected, some of which are also used in our study, such as Kurtosis, Energy, Run Length non-uniformity. The energy-derived features such as Energy and Total Energy are closely related to the intensity of gray levels. These features were found crucial in several studies related to breast cancer, such as the proposed one, and in DCE-MRI, in which a strong correlation is present with the temporal evolution of the contrast agent Prinzi et al. (2024c). These radiomic features describe the change in gray intensity after administration of contrast medium, which is more rapid in malignant lesions, so these latter shows a more rapid increase in energy, viceversa for benign lesions. In breast ultrasound evaluation, excluding completely anechoic lesions, gray levels related to the echogenicity of a focal lesion may vary, but most malignant lesions are hypoechoic Rahbar et al. (1999). Although the correlation between gray level intensity and the malignant lesion is different in DCE-MRI than in ultrasonographic features, even in ultrasonographic radiomic analysis, we can assert that both Energy and Total Energy emerge as characterizing features in ultrasonography, thereby laying a solid foundation for subsequent clinical investigations. The significance of signal intensity in ultrasound has been demonstrated in various anatomical districts Papini et al. (2002); Tessler et al. (2017), indicating its potential importance for breast imaging as well. Another example is the correlation between lesion heterogeneity, habitats’ presence, and malignancy Marusyk and Polyak (2010); Diaz-Cano (2012). The entropy is a quantity that measures the degree of disorder in a system. Among radiomic features, entropy is directly proportional to the heterogeneity of the lesion ’system’ (e.g. due to prominent vascularization and presence of different types of cells) and thus its malignancy.
As a consequence, Rad4XCNN, by establishing a link between deep features and related radiomic features that can be interpreted from a radiological point of view, can reduce the gap between black-box models and the explainability properties required by radiologists. Our results confirm an interesting overlap between the radiomics features and radiologist insights. This enables deep features to be considered clinical biomarkers such as radiomic features, with the advantage of obtaining more performing models. The missing link in this chain - the correlation between deep features and radiomics - is provided by Rad4XCNN. To provide more detail on this last aspect, Figure 5 depicts the correlation trend.
5.5 Clinical Implications of Rad4XCNN Use
Rad4XCNN contributes to improving current clinical practices in the radiologic domain. In particular, three main aspects were addressed:
  • model clinical validation through quantitative radiomic features: Rad4XCNN stands out for its focus on clinical validation, using quantitative radiomic variables that provide an objective and measurable view of image global characteristics, allowing for more accurate and reproducible radiologist-based assessment. This aspect is essential for model reliability;
  • reduction of inter-operator variability in interpreting model findings: Rad4XCNN improves model explanation, overcoming the qualitative interpretation of saliency maps, which may be subject to inter-operator variability. As a result, Rad4XCNN enhances clinicians’ trust and facilitates the adoption of predictive models into everyday workflows;
  • integration of intelligible features within the clinical practice: The integration of radiomic features into the interpretation of medical images provides an added layer of confidence in results validation. Radiologists can now combine traditional image analysis with the richness of radiomic features. This fusion enhances clinical decision-making and enhances the traditional reporting process with new possible features to use as biomarkers.
5.6 Rad4XCNN Limitations
From a methodological perspective, Rad4XCNN application requires ROIs segmentation for radiomic feature extraction. In contrast, deep learning models typically automatically learn the most important features (the ROI is not needed). In particular, for Rad4XCNN, it is crucial to have pre-defined segmentation of ROIs to ensure that radiomic features are extracted solely from pathological areas, avoiding irrelevant information from surrounding healthy tissue. However, although segmentation is a resource-intensive process, several tools are available nowadays for accurate and rapid segmentation. From a clinical perspective, it is essential for radiologists to understand the meaning of each radiomic feature to derive clinical insights accurately. While the number of extracted features is often substantial, radiomic features have become increasingly standardized and are now being widely incorporated into clinical literature. Furthermore, feature extraction software is typically accompanied by detailed documentation, making it easier for non-experts to comprehend and utilize these features effectively.